A wave is propagated in a waveguide at frequency of 9 GHz and separation is 2 cm between walls calculate the phase velocity for dominant mode.


1
Expert's answer
2021-01-11T11:59:40-0500

Answer

Relationship between group and phase velocity

"v_pv_g=c^2"

So phase velocity is given by

"v_p=\frac{c}{\sqrt{1-(\frac{f_c}{f}) ^2}}"

Putting all given value in question

"c=3\times10^8 m/s"

And all other

"v_p=5\times10^8m/s"